### v2

# Scattering problem for Zakharov-Kuznetsov equation in three space dimensions

Abstract.
This paper is a continuation of our previous study [25] on the scattering problem for the Zakharov-Kuznetsov equation (ZK). When the space dimension is three, we construct a global solution to (ZK) which scatters to a given free solution without smallness assumption on the asymptotic states.
Key words and phrases:
Zakharov-Kuznetsov equation, scattering
2010 Mathematics Subject Classification:
Primary 35Q53, Secondary 35P25
1. Introduction
This paper is a continuation of our previous study [25] on the scattering problem for the Zakharov-Kuznetsov equation. In this paper we focus on the Zakharov-Kuznetsov equation in three dimensions:
(1.1) $\partial_{t}u+\partial_{x_{1}}\Delta u=\partial_{x_{1}}(u^{2}),\qquad(t,x)\in\mathbb{R}\times\mathbb{R}^{3},$
where $u:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}$ is an unknown function, $x=(x_{1},x_{2},x_{3})$ and $\Delta$ is Laplacian on $\mathbb{R}^{3}$. Equation (1.1) was derived by Zakharov-Kuznetsov [27] to describe unidirectional wave propagation in a magnetized plasma. Note that Laedke-Spatschek [18] derived (1.1) from the basic hydrodynamical equations. Furthermore, Lannes-Linares-Saut [19] gave the rigorous justification of (1.1) from the Euler-Poisson system for a uniformly magnetized media.
Equation (1.1) has the conservation of mass : for any $t\in\mathbb{R}$,
(1.2) $M[u](t):=\frac{1}{2}\int_{\mathbb{R}^{3}}u(t,x)^{2}dx=M[u](0),$
and the conservation of energy : for any $t\in\mathbb{R}$,
(1.3) $\ \ \ E[u](t):=\frac{1}{2}\int_{\mathbb{R}^{3}}|(\nabla u)(t,x)|^{2}dx-\frac{1}{6}\int_{\mathbb{R}^{3}}u(t,x)^{3}dxdy=E[u](0).$
The Zakharov-Kuznetsov equation on $\mathbb{R}^{d}$ :
(1.4) $\partial_{t}u+\partial_{x_{1}}\Delta u=\partial_{x_{1}}(u^{2}),\qquad(t,x)\in\mathbb{R}\times\mathbb{R}^{d}$
has been studied from the point of view of well-posedness [5, 10, 12, 13, 15, 20, 21, 24], and stability of soliton [3, 4, 6, 17, 26] etc. Concerning the scattering problem for (1.4), from the fact that the solution of the linear equation associated with (1.4) decays like $O(t^{-d/2})$ in $L^{\infty}$ as $t\to\infty$ (see [14, Theorem 3.2] for instance), and from the point of view of the linear scattering theory (see [23] for instance), we expect that if $d\geq 3$, then (at least small) solution to (1.4) scatters to the free solution. Herr-Kinoshita [12] proved the small data scattering for the initial value problem of (1.4) with $d\geq 5$ in the scaling critical Sobolev space. Furthermore, they proved the scattering for (1.4) with $d=4$ when the initial data is small and radial in the last (d1)(d-1) variables. For two dimensional case $d=2$, the author [25] proved the existence of small global solutions to (1.4) which scatters to a given free solution. See also Farah-Linares-Pastor [7], Anjolras [1], and Correia-Kinoshita [2] for the scattering results on (1.4) with $d=2$ and power type nonlinearity with degree higher than two.
In this paper we consider the scattering problem for (1.4) with physically important case $d=3$ in the framework of the final state problem. To state the our main theorem, we introduce several notation. For $0<\delta<1$, we define a semi-normed space $(X_{\delta},\|\cdot\|_{X_{\delta}}$) by
(1.5) $X_{\delta}$ $:=$ $\{f\in{{\mathcal{S}}}(\mathbb{R}^{3});\|f\|_{X_{\delta}}<\infty\},$
$\|f\|_{X_{\delta}}$ $:=$ $\||\partial_{x_{1}}|^{-\delta}\langle x\rangle f\|_{W_{x}^{6,1}}+\|(3\partial_{x_{1}}^{2}-\partial_{x_{2}}^{2}-\partial_{x_{3}}^{2})^{-2}\langle x\rangle f\|_{H_{x}^{7}}$
$+\|(3\partial_{x_{1}}^{2}-\partial_{x_{2}}^{2}-\partial_{x_{3}}^{2})^{-3}f\|_{H_{x}^{8}},$
where $x=(x_{1},x_{2},x_{3})$, $\langle x\rangle=\sqrt{1+|x|^{2}}$, and $P(-i\nabla)={{\mathcal{F}}}^{-1}P(\xi){{\mathcal{F}}}$ for $P=|\xi_{1}|^{-\delta}$ and $(-3\xi_{1}^{2}+\xi_{2}^{2}+\xi_{3}^{2})^{-m}$, $m=2,3$. Let $\{V(t)\}_{t\in\mathbb{R}}$ be a unitary group on $L^{2}$ generated by $\partial_{x_{1}}\Delta$. Then we have the following.
Theorem 1.1.
Let $0<\delta<1$. Then for any $u_{+}\in X_{\delta}$, there exists a unique global solution $u\in C(\mathbb{R};H^{1}(\mathbb{R}^{2}))$ to (1.1) satisfying
(1.6) $\|u(t)-V(t)u_{+}\|_{H_{x}^{3}}\lesssim t^{-\alpha}$
for any $t\geq 1$, where $\alpha>1/2$. Similar result holds for negative time direction.
Remark 1.
In [25], we proved the scattering result similar to Theorem 1.1 in two dimensional case under the smallness assumption on the asymptotic states $u_{+}$. Note that in Theorem 1.1, we do not require smallness assumption on the asymptotic states thanks to good time decay of the free solution.
Remark 2.
The differential operator $3\partial_{x_{1}}^{2}-\partial_{x_{2}}^{2}-\partial_{x_{3}}^{2}$ in $X_{\delta}$ appears naturally in study of the linear/nonlinear scattering for (1.1). For example, the solution to the linear equation for (1.1) satisfies the following time decay estimate :
$\||\partial_{x_{1}}|^{\frac{1}{2}}|3\partial_{x_{1}}^{2}-\partial_{x_{2}}^{2}-\partial_{x_{3}}^{2}|^{\frac{1}{2}}V(t)Pf\|_{L_{x}^{\infty}}\lesssim t^{-\frac{3}{2}}\|f\|_{L_{x}^{1}},$
where $P$ is a suitable projection (see [14, Theorem 3.2]).
Remark 3.
In the definition of the function space $X_{\delta}$, we required that $f\in{{\mathcal{S}}}$ for the simplicity of the argument. This requirement can be relaxed by the density argument for example. However, we do not discuss about it in this paper.
We now give the outline of the proof of Theorem 1.1. As in [25], for given final data $u_{+}\in X_{\delta}$, we introduce a new unknown function
$w(t,x):=u(t,x)-u_{1}(t,x)-u_{2}(t,x),$
where $u_{1}$ and $u_{2}$ are given by
(1.7) $u_{1}(t,x)$ $=$ $[V(t)u_{+}](x),$
$u_{2}(t,x)$ $=$ $-\partial_{x_{1}}\int_{t}^{\infty}V(t-\tau)[u_{1}(\tau)^{2}]d\tau$
$=$ $-\partial_{x_{1}}\int_{t}^{\infty}V(t-\tau)[(V(\tau)u_{+})^{2}]d\tau.$
Let us derive the evolution equation for $w$. Let ${{\mathcal{L}}}=\partial_{t}+\partial_{x_{1}}\Delta$. Since ${{\mathcal{L}}}u_{1}=0$ and ${{\mathcal{L}}}u_{2}=\partial_{x_{1}}(u_{1}^{2})$, we have
${{\mathcal{L}}}w={{\mathcal{L}}}u-{{\mathcal{L}}}u_{1}-{{\mathcal{L}}}u_{2}={{\mathcal{L}}}u-\partial_{x_{1}}(u_{1}^{2}).$
If $u$ satisfies (1.1), then
${{\mathcal{L}}}u$ $=$ $\partial_{x_{1}}(u^{2})$
$=$ $\partial_{x_{1}}\{(w+u_{1}+u_{2})^{2}\}$
$=$ $\partial_{x_{1}}\{w^{2}+2(u_{1}+u_{2})w+u_{1}^{2}+2u_{1}u_{2}+u_{2}^{2}\}$
$=$ $\partial_{x_{1}}(w^{2})+2\partial_{x_{1}}\{(u_{1}+u_{2})w\}+\partial_{x_{1}}(u_{1}^{2}+2u_{1}u_{2}+u_{2}^{2}).$
Hence we see that $w$ satisfies
(1.9) $\ \ \ {{\mathcal{L}}}w=\partial_{x_{1}}(w^{2})+2\partial_{x_{1}}\left\{(u_{1}+u_{2})w\right\}+\partial_{x_{1}}(2u_{1}u_{2}+u_{2}^{2}).$
To show Theorem 1.1, we prove the existence of solution $w$ to (1.9) satisfying
$\sup_{t\in[T,\infty)}t^{\alpha}\left(\|w(t)\|_{H_{x}^{3}}+\||\partial_{x_{1}}|^{\frac{\nu}{2}}w(\tau)\|_{L_{\tau}^{\frac{6}{3-2\nu}}(t,\infty;W_{x}^{2,\frac{2}{\nu}})}\right)<\infty$
for suitable $\alpha>0$, $T>0$ and $0<\nu<1/2$. We first note that $w$ and $u_{1}$ can be easily estimated by the energy and the linear dispersive estimates (Strichartz estimates, see Lemma 2.1 below). The main difficulty to prove Theorem 1.1 lies on bilinear dispersive estimates for $u_{2}$. More precisely, we need $L^{2}$ estimate for the bilinear oscillatory integral :
$\ \ \ \ \ u_{2}$ $=$ $-\partial_{x_{1}}\int_{t}^{\infty}V(t-\tau)\left[(V(\tau)u_{+})^{2}\right]d\tau$
$=$ $-{{\mathcal{F}}}_{\xi\mapsto x}^{-1}\left[\xi_{1}e^{it\xi_{1}|\xi|^{2}}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}e^{-i\tau\phi(\xi,\eta)}\widehat{u}_{+}(\xi-\eta)\widehat{u}_{+}(\eta)d\eta d\tau\right],$
where $\xi=(\xi_{1},\xi_{2},\xi_{3})$, $\eta=(\eta_{1},\eta_{2},\eta_{3})$ and $\phi(\xi,\eta)=\xi_{1}|\xi|^{2}-(\xi_{1}-\eta_{1})|\xi-\eta|^{2}-\eta_{1}|\eta|^{2}$. To derive time decay estimates for (1) in $L^{2}$, we employ so called space-time resonance method which is developed by Gustafson-Nakanishi-Tsai [11] and Germain-Masmoudi-Shatah [8, 9] etc. In this paper, we crucially use the “null structure” of the nonlinear term which can be represented as the algebraic identity
(1.11) $\xi_{1}=\left(\sum p(\xi)q(\eta)\right)\phi+\sum_{j=1}^{3}\left(\sum r(\xi)s(\eta)\right)\partial_{\eta_{j}}\phi$
with suitable polynomials $p,r$ and rational functions $q,s$, see Lemma 3.2 below for detail. Combining (1.11) with integration by parts both in $\tau$ and $\eta$, we derive the time decay of (1) in $L^{2}$. Note that the null structure of the nonlinearity is employed in various contexts to study global dynamics of nonlinear PDEs since the pioneering work by Klainerman [16]. We note that this approach was also used in [25] to prove the existence of small global solution to (1.1) with $d=2$ which scatters to the given free solution. Furthermore, in [25], the author transformed (1.4) into the equation which is symmetric with respect to $x_{1}$ and $x_{2}$. Thanks to this transform, the problem became more transparent in two dimensional case. On the other hand, the similar transform is not known for (1.4) with $d\neq 2$. Therefore derivation of the key identity (1.11) for $d=3$ is more complicated compared to two dimensional case. Once we obtain $L^{2}$ estimate for $u_{2}$, we have an existence of solution to (1.9) by the compactness argument.
We introduce several notations and function spaces which are used throughout this paper. For $f\in{{\mathcal{S}}}^{\prime}(\mathbb{R}^{3})$, $\hat{f}(\xi)$ denotes the Fourier transform of $f$. Let $\langle\xi\rangle=\sqrt{|\xi|^{2}+1}$. The differential operator $\langle\nabla\rangle^{s}=(1-\Delta)^{s/2}$ denotes the Bessel potential of order s-s. For $1\leq p,q\leq\infty$, $L_{\tau}^{p}(t,\infty;L_{x}^{q})$ is defined as follows:
$L_{\tau}^{p}(t,\infty;L_{x}^{q})$ $=$ $\{u\in{{\mathcal{S}}}^{\prime}(\mathbb{R}^{1+3});\|u\|_{L_{\tau}^{p}(t,\infty;L_{x}^{q})}<\infty\},$
$\|u\|_{L_{\tau}^{p}(t,\infty;L_{x}^{q})}$ $=$ $\left\|\|u(\tau)\|_{L_{x}^{q}}\right\|_{L_{\tau}^{p}(t,\infty)}.$
We will use the inhomogeneous Sobolev spaces
$W^{s,q}=\{f\in\mathcal{S}^{\prime}(\mathbb{R}^{3});\|f\|_{W^{s,q}}=\|\langle\nabla\rangle^{s}f\|_{L^{q}}<\infty\},$
where $s\in\mathbb{R}$ and $1\leq q\leq\infty$. We denote $H^{s}=W^{s,2}$. We denote $A\lesssim B$ if there exists a constant $C>0$ such that $A\leq CB$ holds and $A\sim B$ if $A\lesssim B\lesssim A$.
The outline of the paper is as follows. In Section 2, we give the decay and Strichartz estimates for the linearized equation of (1.9). In Section 3, we derive the key bilinear dispersive estimates. Finally, in Section 4, we prove Theorem 1.1. In Appendix, we give the proof of Lemma 3.2.
2. Linear Dispersive Estimates
In this section we derive the linear estimates associated with (1.9):
(2.1) $\left\{\begin{array}[]{l}\partial_{t}w+\partial_{x_{1}}\Delta w=0,\qquad\quad(t,x)\in\mathbb{R}\times\mathbb{R}^{3},\\ w(0,x)=f(x),\qquad\qquad\ x\in\mathbb{R}^{3}.\end{array}\right$.
Let us recall that $V(t)=e^{-t\partial_{x_{1}}\Delta}$ is the unitary group on $L^{2}$ generated by $\partial_{x_{1}}\Delta$. Then, the solution to (2.1) can be written as $V(t)f$.
We have the following decay and the Strichartz estimates for (2.1).
Lemma 2.1.
(i) Let $0<a<1$ and $0\leq b\leq 1$. Then for any $t>0$, we have
(2.2) $\||\partial_{x_{1}}|^{ab}V(t)f\|_{L_{x}^{q}}$ $\lesssim$ $t^{-b(1+\frac{a}{3})}\|f\|_{L_{x}^{q^{\prime}}},$
where q=2/(1b)q=2/(1-b) and $q^{\prime}$ is the Hölder conjugate exponent of $q$.
(ii) Let $0<a<1$ and $0<b<(1+a/3)^{-1}$. Then, we have
(2.3) $\left\||\partial_{x_{1}}|^{ab}\int_{\tau}^{\infty}V(\tau-\tau^{\prime})F(\tau^{\prime})d\tau^{\prime}\right\|_{L_{\tau}^{p}(t,\infty;L_{x}^{q})}\lesssim\|F\|_{L_{\tau}^{p^{\prime}}(t,\infty;L_{x}^{q^{\prime}})},$
where $p=2/\{b(1+a/3)\}$ and q=2/(1b)q=2/(1-b).
Proof of Lemma 2.1..
See [20, Lemma 3.3] for the proof of (2.2), and [20, Proposition 3.1] for the proof of (2.3). ∎
3. Bilinear Dispersive Estimates
In this section we derive $L^{2}$ estimate for $u_{2}$ defined by (1) which is key to prove Theorem 1.1. We show the following.
Proposition 3.1.
Let $0<\delta<1$. Then for any $t>0$, we have
(3.1) $\ \ \ \ \ \left\|\partial_{x_{1}}\int_{t}^{\infty}V(t-\tau)\left[(V(\tau)f)(V(\tau)g)\right]d\tau\right\|_{L_{x}^{2}}\lesssim t^{-1-\frac{\delta}{3}}\|f\|_{Y_{\delta}}\|g\|_{Y_{\delta}},$
where
$Y_{\delta}$ $=$ $\{f\in{{\mathcal{S}}}(\mathbb{R}^{3});\|f\|_{Y_{\delta}}<\infty\},$
$\|f\|_{Y_{\delta}}$ $=$ $\||\partial_{x_{1}}|^{-\delta}\langle x\rangle f\|_{W_{x}^{2,1}}+\|(3\partial_{x_{1}}^{2}-\partial_{x_{2}}^{2}-\partial_{x_{3}}^{2})^{-2}\langle x\rangle f\|_{H_{x}^{3}}$
$+\|(3\partial_{x_{1}}^{2}-\partial_{x_{2}}^{2}-\partial_{x_{3}}^{2})^{-3}f\|_{H_{x}^{4}}.$
To prove Proposition 3.1, we need to do some preparation. Simple calculation yields
$\partial_{x_{1}}\int_{t}^{\infty}V(t-\tau)\left[(V(\tau)f)(V(\tau)g)\right]d\tau$
$=$ ${{\mathcal{F}}}_{\xi\mapsto x}^{-1}\left[\xi_{1}e^{it\xi_{1}|\xi|^{2}}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}e^{-i\tau\phi(\xi,\eta)}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau\right](x),$
where $\xi=(\xi_{1},\xi_{2},\xi_{3}),\eta=(\eta_{1},\eta_{2},\eta_{3})$ and the resonant function $\phi$ is given by
(3.3) $\phi(\xi,\eta)$ $=$ $\xi_{1}|\xi|^{2}-(\xi_{1}-\eta_{1})|\xi-\eta|^{2}-\eta_{1}|\eta|^{2}.$
Therefore, to prove Proposition 3.1, we need to estimate
(3.4) $I(f,g):=\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}e^{-i\tau\phi(\xi,\eta)}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau.$
We evaluate (3.4) by using the space-time resonance method. To this end, we derive the following key algebraic identity.
Lemma 3.2.
We have
(3.5) $\xi_{1}=\psi_{\text{time}}(\xi,\eta)+\psi_{\text{space}}(\xi,\eta),$
where $\psi_{\text{time}}$ and $\psi_{\text{space}}$ are given by
(3.6) $\psi_{\text{time}}(\xi,\eta)$ $=$ $\left(A_{0}(\eta)+\xi_{1}B_{0,1}(\eta)+\xi_{2}B_{0,2}(\eta)+\xi_{3}B_{0,3}(\eta)\right)\phi$
$\ \ \ \psi_{\text{space}}(\xi,\eta)$ $=$ $\left(A_{1}(\eta)+\xi_{1}B_{1,1}(\eta)+\xi_{2}B_{1,2}(\eta)+\xi_{3}B_{1,3}(\eta)\right)\partial_{\eta_{1}}\phi$
$+\left(A_{2}(\eta)+\xi_{1}B_{2,1}(\eta)+\xi_{2}B_{2,2}(\eta)\right$.
$\qquad\qquad\qquad\qquad\left.+\xi_{2}\xi_{3}C_{2,1}(\eta)+\xi_{3}^{2}C_{2,2}(\eta)\right)\partial_{\eta_{2}}\phi$
$+\left(A_{3}(\eta)+\xi_{1}B_{3,1}(\eta)+\xi_{3}B_{3,2}(\eta)\right$.
$\qquad\qquad\qquad\qquad\left.+\xi_{2}\xi_{3}C_{3,1}(\eta)+\xi_{2}^{2}C_{3,2}(\eta)\right)\partial_{\eta_{3}}\phi,$
where $A_{j}$, $B_{j,k}$ and $C_{j,k}$ are rational functions in $\eta$ satisfying
$|A_{0}(\eta)|$ $\lesssim$ $(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2},$
$|\partial_{\eta_{j}}^{i}A_{j}(\eta)|$ $\lesssim$ $\left\{\begin{aligned} &(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{\frac{3}{2}}(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\qquad(i=0),\\ &(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\\ &+(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{2}|3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2}|^{-3}\qquad(i=1)\end{aligned}\right$.
for $j=1,2,3$,
$|B_{0,k}(\eta)|$ $\lesssim$ $(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}$
for $k=1,2,3$,
$|\partial_{\eta_{j}}^{i}B_{j,k}(\eta)|$ $\lesssim$ $\left\{\begin{aligned} &(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\quad(i=0),\\ &(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{\frac{1}{2}}(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\\ &+(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{\frac{3}{2}}|3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2}|^{-3}\quad(i=1)\end{aligned}\right$.
for (j,k)=(1,1),(1,2),(1,3),(2,1),(2,2),(3,1),(3,2)(j,k)=(1,1),(1,2),(1,3),(2,1),(2,2),(3,1),(3,2),
$|\partial_{\eta_{j}}^{i}C_{j,k}(\eta)|$ $\lesssim$ $\left\{\begin{aligned} &(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{\frac{1}{2}}(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\quad(i=0),\\ &(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\\ &+(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})|3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2}|^{-3}\quad(i=1),\end{aligned}\right$.
for $(j,k)=(2,1),(2,2),(3,1),(3,2)$.
We shall prove Lemma 3.2 in Appendix.
Proof of Proposition 3.1..
By using (3.5), we split $I(f,g)$ defined by (3.4) into the following two terms:
$I(f,g)$ $=$ $\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\psi_{\text{time}}(\xi,\eta)e^{-i\tau\phi(\xi,\eta)}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$+\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\psi_{\text{space}}(\xi,\eta)e^{-i\tau\phi(\xi,\eta)}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$=:$ $I_{\text{time}}(f,g)+I_{\text{space}}(f,g),$
where $\psi_{\text{time}}$ and $\psi_{\text{space}}$ are given by (3.6) and (3.2), respectively.
We first evaluate $I_{\text{time}}(f,g)$. We treat the second term of $\psi_{\text{time}}$, i.e., $\xi_{1}B_{0,1}(\eta)\phi$ only since the other terms can be treated in a similar way. Let
$I_{\text{time},2}(f,g):=\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}B_{0,1}(\eta)\phi(\xi,\eta)e^{-i\tau\phi(\xi,\eta)}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau.$
Integrating in $\tau$, we have
$I_{\text{time},2}(f,g)$ $=$ $i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\partial_{\tau}\left\{e^{-i\tau\phi(\xi,\eta)}\right\}B_{0,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$=$ $i\limsup_{T\to\infty}\xi_{1}\int_{\mathbb{R}^{3}}e^{-iT\phi(\xi,\eta)}B_{0,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta$
$-i\xi_{1}\int_{\mathbb{R}^{3}}e^{-it\phi(\xi,\eta)}B_{0,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta.$
Hence, by the Plancherel theorem and noting $\xi_{1}=(\xi_{1}-\eta_{1})+\eta_{1}$, we obtain
$\left\|I_{\text{time},2}(f,g)\right\|_{L_{\xi}^{2}}$
$\lesssim$ $\limsup_{T\to\infty}\left\|\xi_{1}\int_{\mathbb{R}^{3}}{{\mathcal{F}}}[V(T)f](\xi-\eta){{\mathcal{F}}}\left[B_{0,1}(-i\nabla)V(T)g\right](\eta)d\eta\right\|_{L_{\xi}^{2}}$
$+\left\|\xi_{1}\int_{\mathbb{R}^{3}}{{\mathcal{F}}}[V(t)f](\xi-\eta){{\mathcal{F}}}\left[B_{0,1}(-i\nabla)V(t)g\right](\eta)d\eta\right\|_{L_{\xi}^{2}}$
$\lesssim$ $\limsup_{T\to\infty}\|\partial_{x_{1}}V(T)f\|_{L_{x}^{\infty}}\left\|B_{0,1}(-i\nabla)V(T)g\right\|_{L_{x}^{2}}$
$+\limsup_{T\to\infty}\|V(T)f\|_{L_{x}^{\infty}}\left\|\partial_{x_{1}}B_{0,1}(-i\nabla)V(T)g\right\|_{L_{x}^{2}}$
$+\|\partial_{x_{1}}V(t)f\|_{L_{x}^{\infty}}\left\|B_{0,1}(-i\nabla)V(t)g\right\|_{L_{x}^{2}}$
$+\|V(t)f\|_{L_{x}^{\infty}}\left\|\partial_{x_{1}}B_{0,1}(-i\nabla)V(t)g\right\|_{L_{x}^{2}}.$
By the decay estimate (Lemma 2.1 (2.2)) and the inequality for $B_{0,1}$ in Lemma 3.2, we have
$\left\|I_{\text{time},2}(f,g)\right\|_{L_{\xi}^{2}}\lesssim t^{-1-\frac{\delta}{3}}\|f\|_{Y_{\delta}}\|g\|_{Y_{\delta}}.$
In a similar way we have
(3.9) $\left\|I_{\text{time}}(f,g)\right\|_{L_{\xi}^{2}}\lesssim t^{-1-\frac{\delta}{3}}\|f\|_{Y_{\delta}}\|g\|_{Y_{\delta}}.$
Next we evaluate $I_{\text{space}}(f,g)$. We treat the second term of $\psi_{\text{space}}$, i.e., $\xi_{1}B_{1,1}(\eta)\partial_{\eta_{1}}\phi$ only since the other terms can be treated in a similar way. Let
$I_{\text{space,2}}(f,g):=\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}B_{1,1}(\eta)\partial_{\eta_{1}}\phi(\xi,\eta)e^{-i\tau\phi(\xi,\eta)}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau.$
By an integration by parts in $\eta_{1}$, we have
$I_{\text{space,2}}(f,g)$ $=$ $i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}\partial_{\eta_{1}}\left\{e^{-i\tau\phi(\xi,\eta)}\right\}B_{1,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$=$ $-i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}e^{-i\tau\phi(\xi,\eta)}\partial_{\eta_{1}}B_{1,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$+i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}e^{-i\tau\phi(\xi,\eta)}B_{1,1}(\eta)\partial_{\eta_{1}}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$-i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}e^{-i\tau\phi(\xi,\eta)}B_{1,1}(\eta)\widehat{f}(\xi-\eta)\partial_{\eta_{1}}\widehat{g}(\eta)d\eta d\tau.$
Hence, by the Plancherel theorem and noting $\xi_{1}=(\xi_{1}-\eta_{1})+\eta_{1}$, we obtain
$\left\|I_{\text{space},2}(f,g)\right\|_{L_{\xi}^{2}}$
$\lesssim$ $\biggl\|\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}{{\mathcal{F}}}[V(t)f](\xi-\eta){{\mathcal{F}}}\left[(\partial_{\eta_{1}}B_{1,1})(-i\nabla)V(t)g\right](\eta)d\eta d\tau\biggl\|_{L_{\xi}^{2}}$
$+\left\|\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}{{\mathcal{F}}}[V(t)x_{1}f](\xi-\eta){{\mathcal{F}}}\left[B_{1,1}(-i\nabla)V(t)g\right](\eta)d\eta d\tau\right\|_{L_{\xi}^{2}}$
$+\left\|\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}{{\mathcal{F}}}[V(t)f](\xi-\eta){{\mathcal{F}}}\left[B_{1,1}(-i\nabla)V(t)x_{1}g\right](\eta)d\eta d\tau\right\|_{L_{\xi}^{2}}.$
By the decay estimate (Lemma 2.1 (2.2)) and the inequality for $B_{1,1}$ in Lemma 3.2, we have
$\left\|I_{\text{space},2}(f,g)\right\|_{L_{\xi}^{2}}\lesssim t^{-1-\frac{\delta}{3}}\|f\|_{Y_{\delta}}\|g\|_{Y_{\delta}}.$
In a similar way we have
(3.10) $\left\|I_{\text{space}}(f,g)\right\|_{L_{\xi}^{2}}\lesssim t^{-1-\frac{\delta}{3}}\|f\|_{Y_{\delta}}\|g\|_{Y_{\delta}}.$
Combining (3), (3.4), (3), (3.9) and (3.10), we have
$\left\|\partial_{x_{1}}\int_{t}^{\infty}V(t-\tau)\left[(V(\tau)f)(V(\tau)g)\right]d\tau\right\|_{L_{x}^{2}}$
$=$ $\left\|I(f,g)\right\|_{L_{x}^{2}}$
$\lesssim$ $\left\|I_{\text{time}}(f,g)\right\|_{L_{\xi}^{2}}+\left\|I_{\text{space}}(f,g)\right\|_{L_{\xi}^{2}}$
$\lesssim$ $t^{-1-\frac{\delta}{3}}\|f\|_{Y_{\delta}}\|g\|_{Y_{\delta}}.$
Hence we obtain (3.1). This completes the proof of Proposition 3.1. ∎
Lemma $3.$
Let $0<\delta<1$ and let $u_{1}$ and $u_{2}$ be given by (1.7) and (1). Then for any $t>0$, we have
(3.11) $\|u_{1}(t)\|_{W_{x}^{4,\infty}}$ $\lesssim$ $t^{-1-\frac{\delta}{3}}\|u_{+}\|_{X_{\delta}},$
(3.12) $\|u_{2}(t)\|_{H_{x}^{4}}$ $\lesssim$ $t^{-1-\frac{\delta}{3}}\|u_{+}\|_{X_{\delta}}^{2},$
where the semi-norm $\|\cdot\|_{X_{\delta}}$ is given by (1.5).
Proof of Lemma $3.$
By Lemma 2.1 (2.2), we have
$\|u_{1}(t)\|_{W_{x}^{4,\infty}}=\|V(t)u_{+}\|_{W_{x}^{4,\infty}}\lesssim t^{-1-\frac{\delta}{3}}\||\partial_{x_{1}}|^{-\delta}u_{+}\|_{W_{x}^{4,1}},$
which yields (3.11).
To show (3.12), we note
$\|u_{2}(t)\|_{H_{x}^{4}}\sim\|u_{2}(t)\|_{L_{x}^{2}}+\|\Delta^{2}u_{2}(t)\|_{L_{x}^{2}}.$
We define
$B(f,g):=\partial_{x_{1}}\int_{t}^{\infty}V(t-\tau)\left[(V(\tau)f)(V(\tau)g)\right]d\tau.$
Since $\xi=(\xi-\eta)+\eta$, we see
$\nabla B(f,g)=B(\nabla f,g)+B(f,\nabla g).$
In a similar way, using the identity $|\xi|^{2}=|\xi-\eta|^{2}+2(\xi-\eta)\cdot\eta+|\eta|^{2}$, we have
$\Delta B(f,g)=B(\Delta f,g)+2\sum_{j=1}^{3}B(\partial_{x_{j}}f,\partial_{x_{j}}g)+B(f,\Delta g).$
Therefore
$\Delta^{2}B(f,g)$ $=$ $\Delta B(\Delta f,g)+2\sum_{j=1}^{3}\Delta B(\partial_{x_{j}}f,\partial_{x_{j}}g)+\Delta B(f,\Delta g)$
$=$ $B(\Delta^{2}f,g)+4\sum_{j=1}^{3}B(\partial_{x_{j}}\Delta f,\partial_{x_{j}}g)$
$+2B(\Delta f,\Delta g)+4\sum_{j=1}^{3}\sum_{k=1}^{3}B(\partial_{x_{j}}\partial_{x_{k}}f,\partial_{x_{j}}\partial_{x_{k}}g)$
$+4\sum_{j=1}^{3}B(\partial_{x_{j}}f,\partial_{x_{j}}\Delta g)+B(f,\Delta^{2}g).$
Combining Proposition 3.1 (3.1) with the above identities, we have (3.12). ∎
4. Proof of Theorem 1.1
In this section we complete the proof of Theorem 1.1. To prove Theorem 1.1, we show the existence of solution $w$ to (1.9) with $w\to 0$ in $H^{3}(\mathbb{R}^{3})$ as $t\to\infty$.
Let
$N(w,u_{1},u_{2})$ $:=$ $\partial_{x_{1}}(w^{2})+2\partial_{x_{1}}\left\{(u_{1}+u_{2})w\right\}+\partial_{x_{1}}(2u_{1}u_{2}+u_{2}^{2}).$
Then (1.9) can be rewritten as
(4.1) $\partial_{t}w+\partial_{x_{1}}\Delta w=N(w,u_{1},u_{2}).$
To show the existence of solution $w$ to (1.9) with $w\to 0$ in $H^{3}(\mathbb{R}^{3})$, we consider the regularized equation associated with (4.1) :
$\partial_{t}w_{\lambda,\mu}+\partial_{x_{1}}\Delta w_{\lambda,\mu}$
$=$ $(1+\lambda t)^{-5}\rho_{\mu}\ast N(\rho_{\mu}\ast w,\rho_{\mu}\ast u_{1},\rho_{\mu}\ast u_{2}),$
where $0<\lambda,\mu<1$, $\rho\in C_{0}^{\infty}(\mathbb{R}^{3})$ satisfies $\rho\geq 0$ and $\int\rho(x)dx=1$, and $\rho_{\mu}(x)=\mu^{-3}\rho(x/\mu)$.
Thanks to the regularizing factor $\rho_{\mu}\ast$ and the time decaying factor $(1+\lambda t)^{-5}$, by using the contraction mapping principle, we easily see that for any $0<\lambda<1$ and $0<\mu<1$, there exists a $T_{\lambda,\mu}>0$ such that (4) has a unique solution $w_{\lambda,\mu}$ satisfying
$w_{\lambda,\mu}\in\bigcap_{j=1}^{\infty}C^{1}([T_{\lambda,\mu},\infty),H_{x}^{j}),$
$\sup_{t\geq T_{\lambda,\mu}}(1+\lambda t)^{4}\sum_{3i+j\leq 3}\|\partial_{t}^{i}\nabla_{x}^{j}w_{\lambda,\mu}(t)\|_{L_{x}^{2}}<+\infty.$
Again using the regularizing and time decaying factors, the above solution $w_{\lambda,\mu}$ can be extend to $0,\infty)$ without the smallness assumption on $u_{+}$.
We next derive an a priori estimates for $w_{\lambda,\mu}$ independent of $\lambda$ and $\mu$ under the assumption that $r:=\|u_{+}\|_{X_{\delta}}<\infty$, where $\|\cdot\|_{X_{\delta}}$ is defined by (1.5). We abbreviate $w_{\lambda,\mu}$ to $w$. Let
$\|w\|_{Z_{T}}:=\sup_{t\in[T,\infty)}t^{\alpha}\left(\|w(t)\|_{H_{x}^{3}}+\||\partial_{x_{1}}|^{\frac{\nu}{2}}w(\tau)\|_{L_{\tau}^{\frac{6}{3-2\nu}}(t,\infty;W_{x}^{2,\frac{2}{\nu}})}\right),$
where $\alpha>0$, $T>0$ and $0<\nu<1/2$ are fixed later.
We first derive the estimates for $w$ in $H_{x}^{3}$.
Lemma 4.1.
Let $w$ be a solution to (4). Then we have
$\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}$
$\lesssim$ $(r^{3}+r^{4})T^{\alpha-\frac{2}{3}\delta-1}+(r+r^{4})T^{-\frac{\delta}{3}}\|w\|_{Z_{T}}+T^{-\alpha+\frac{\nu}{3}+\frac{1}{2}}\|w\|_{Z_{T}}^{2}$
$+(r+r^{2})T^{-\alpha-\frac{\delta}{3}}\|w\|_{Z_{T}}^{2}+T^{-2\alpha+1}\|w\|_{Z_{T}}^{3}$
$+(r^{3}+r^{4})T^{-11\alpha-\frac{2}{3}\delta-1}\|w\|_{Z_{T}}^{12}+(r+r^{2})T^{-12\alpha-\frac{\delta}{3}}\|w\|_{Z_{T}}^{13},$
where the implicit constants are independent of $\lambda$ and $\mu$.
Proof of Lemma 4.1..
The proof is based on the energy method. Although we derive (4.1) for smooth solution to (4.1), the proof for (4.1) below works for (4).
Let $w$ be a smooth solution to (4.1). Taking the inner product in $L_{x}^{2}$ between (4.1) and $w$ and integrating by parts, we have
$\frac{1}{2}\frac{d}{dt}\|w\|_{L_{x}^{2}}^{2}$ $=$ $\int_{\mathbb{R}^{3}}\partial_{x_{1}}(w^{2})wdx+2\int_{\mathbb{R}^{3}}\partial_{x_{1}}\{(u_{1}+u_{2})w\}wdx$
$+\int_{\mathbb{R}^{3}}\partial_{x_{1}}(2u_{1}u_{2}+u_{2}^{2})wdx$
$=$ $-\int_{\mathbb{R}^{3}}w^{2}\partial_{x_{1}}wdx-2\int_{\mathbb{R}^{3}}(u_{1}+u_{2})w\partial_{x_{1}}wdx$
$+\int_{\mathbb{R}^{3}}\partial_{x_{1}}(2u_{1}u_{2}+u_{2}^{2})wdx$
$=$ $\int_{\mathbb{R}^{3}}\partial_{x_{1}}(u_{1}+u_{2})w^{2}dx+\int_{\mathbb{R}^{3}}\partial_{x_{1}}(2u_{1}u_{2}+u_{2}^{2})wdx$
$\lesssim$ $(\|u_{1}\|_{W_{x}^{1,\infty}}+\|u_{2}\|_{H_{x}^{3}})\|w\|_{L_{x}^{2}}^{2}$
$+(\|u_{1}\|_{W_{x}^{1,\infty}}+\|u_{2}\|_{H_{x}^{3}})\|u_{2}\|_{H_{x}^{3}}\|w\|_{L_{x}^{2}}.$
Hence
$\frac{d}{dt}\|w\|_{L_{x}^{2}}^{14}$ $=$ $7\|w\|_{L_{x}^{2}}^{12}\cdot\frac{d}{dt}\|w\|_{L_{x}^{2}}^{2}$
$\lesssim$ $(\|u_{1}\|_{W_{x}^{1,\infty}}+\|u_{2}\|_{H_{x}^{3}})\|w\|_{L_{x}^{2}}^{14}$
$+(\|u_{1}\|_{W_{x}^{1,\infty}}+\|u_{2}\|_{H_{x}^{3}})\|u_{2}\|_{H_{x}^{3}}\|w\|_{L_{x}^{2}}^{13}.$
Applying $\nabla\Delta$ to (4.1) and taking the inner product in $L_{x}^{2}$ between the resulting equation and $\nabla\Delta w$, we obtain
$\frac{1}{2}\frac{d}{dt}\|\nabla\Delta w\|_{L_{x}^{2}}^{2}$ $=$ $\int_{\mathbb{R}^{3}}\nabla\Delta\partial_{x_{1}}(w^{2})\cdot\nabla\Delta wdx$
$+2\int_{\mathbb{R}^{3}}\nabla\Delta\partial_{x_{1}}\{(u_{1}+u_{2})w\}\cdot\nabla\Delta wdx$
$+\int_{\mathbb{R}^{3}}\nabla\Delta\partial_{x_{1}}(2u_{1}u_{2}+u_{2}^{2})\cdot\nabla\Delta wdx$
$=:$ $I_{1}+I_{2}+I_{3}.$
For $I_{1}$, by integration by parts, we have
$I_{1}$ $=$ $-\int_{\mathbb{R}^{3}}\nabla\Delta(w^{2})\cdot\partial_{x_{1}}\nabla\Delta wdx$
$=$ $-\sum_{j=1}^{3}\int_{\mathbb{R}^{3}}\partial_{x_{j}}\Delta(w^{2})\cdot\partial_{x_{1}}\partial_{x_{j}}\Delta wdx$
$=$ $-4\sum_{j=1}^{3}\int_{\mathbb{R}^{3}}\nabla w\cdot\partial_{x_{j}}\nabla w\partial_{x_{1}}\partial_{x_{j}}\Delta wdx-2\sum_{j=1}^{3}\int_{\mathbb{R}^{3}}\partial_{x_{j}}w\Delta w\partial_{x_{1}}\partial_{x_{j}}\Delta wdx$
$-2\sum_{j=1}^{3}\int_{\mathbb{R}^{3}}w\partial_{x_{j}}\Delta w\partial_{x_{1}}\partial_{x_{j}}\Delta wdx.$
By integration by parts again, we obtain
$I_{1}$ $=$ $4\sum_{j=1}^{3}\int_{\mathbb{R}^{3}}|\partial_{x_{j}}\nabla w|^{2}\partial_{x_{1}}\Delta wdx+4\int_{\mathbb{R}^{3}}\nabla w\cdot\nabla\Delta w\partial_{x_{1}}\Delta wdx$
$+2\int_{\mathbb{R}^{3}}\nabla w\cdot\nabla\Delta w\partial_{x_{1}}\Delta wdx+\int_{\mathbb{R}^{3}}\partial_{x_{1}}w|\nabla\Delta w|^{2}dx$
$=$ $6\int_{\mathbb{R}^{3}}\nabla w\cdot\nabla\Delta w\partial_{x_{1}}\Delta wdx+\int_{\mathbb{R}^{3}}\partial_{x_{1}}w|\nabla\Delta w|^{2}dx+R_{1},$
where
$R_{1}$ $=$ $4\sum_{j=1}^{3}\int_{\mathbb{R}^{3}}|\partial_{x_{j}}\nabla w|^{2}\partial_{x_{1}}\Delta wdx$
$=$ $-8\sum_{j,k=1}^{3}\int_{\mathbb{R}^{3}}\partial_{x_{j}}\nabla w\cdot\partial_{x_{j}}\partial_{x_{k}}\nabla w\partial_{x_{1}}\partial_{x_{k}}wdx$
$\lesssim$ $\|\partial_{x_{1}}w\|_{W_{x}^{1,\frac{2}{\nu}}}\|w\|_{H_{x}^{3}}^{2}.$
In a similar way, we see
(4.9) $|I_{2}|$ $\lesssim$ $(\|u_{1}\|_{W_{x}^{4,\infty}}+\|u_{2}\|_{H_{x}^{4}})\|w\|_{H_{x}^{3}}^{2},$
(4.10) $|I_{3}|$ $\lesssim$ $(\|u_{1}\|_{W_{x}^{4,\infty}}+\|u_{2}\|_{H_{x}^{4}})\|u_{2}\|_{H_{x}^{4}}\|w\|_{H_{x}^{3}}.$
On the other hand
$\frac{d}{dt}\int_{\mathbb{R}^{3}}w(\Delta w)^{2}dx$ $=$ $\int_{\mathbb{R}^{3}}\partial_{t}w(\Delta w)^{2}dx+2\int_{\mathbb{R}^{3}}w\Delta w\partial_{t}\Delta wdx$
$=$ $-2\int_{\mathbb{R}^{3}}w\Delta w\partial_{x_{1}}\Delta^{2}wdx+R_{2},$
where
$R_{2}$ $=$ $-\int_{\mathbb{R}^{3}}\partial_{x_{1}}\Delta w(\Delta w)^{2}dx+\int_{\mathbb{R}^{3}}N(w,u_{1},u_{2})(\Delta w)^{2}dx$
$+2\int_{\mathbb{R}^{3}}w\Delta w\Delta N(w,u_{1},u_{2})dx$
$\lesssim$ $\|w\|_{H_{x}^{3}}^{4}+(\|u_{1}\|_{W_{x}^{3,\infty}}+\|u_{2}\|_{H_{x}^{3}})\|w\|_{H_{x}^{3}}^{3}$
$+(\|u_{1}\|_{W_{x}^{3,\infty}}+\|u_{2}\|_{H_{x}^{3}})\|u_{2}\|_{H_{x}^{3}}\|w\|_{H_{x}^{3}}^{2}.$
By integration by parts, we obtain
$\frac{d}{dt}\int_{\mathbb{R}^{3}}w(\Delta w)^{2}dx$
$=$ $2\int_{\mathbb{R}^{3}}\nabla w\Delta w\cdot\partial_{x_{1}}\nabla\Delta wdx+2\int_{\mathbb{R}^{3}}w\nabla\Delta w\cdot\partial_{x_{1}}\nabla\Delta wdx$
$+R_{2}$
$=$ $-2\int_{\mathbb{R}^{3}}\nabla w\cdot\nabla\Delta w\partial_{x_{1}}\Delta wdx-\int_{\mathbb{R}^{3}}\partial_{x_{1}}w|\nabla\Delta w|^{2}dx+R_{2}.$
Hence, from (4), (4) and (4), we see that for any $M>0$,
$\frac{d}{dt}\left(\|w(t)\|_{H_{x}^{3}}^{2}+6\int_{\mathbb{R}^{3}}w(\Delta w)^{2}dx+M\|w\|_{L_{x}^{2}}^{14}\right)$
$=$ $-4\int_{\mathbb{R}^{3}}\partial_{x_{1}}w|\nabla\Delta w|^{2}dx+\frac{d}{dt}\|w\|_{L_{x}^{2}}^{2}+M\frac{d}{dt}\|w\|_{L_{x}^{2}}^{14}$
$+2I_{2}+2I_{3}+2R_{1}+R_{2}.$
By (4), (4), (4), (4.9), (4.10) and (4), we have
$\|w(t)\|_{H_{x}^{3}}^{2}+6\int_{\mathbb{R}^{3}}w(\Delta w)^{2}dx+M\|w\|_{L_{x}^{2}}^{14}$
$\lesssim$ $\int_{t}^{+\infty}(\|u_{1}(\tau)\|_{W_{x}^{4,\infty}}+\|u_{2}(\tau)\|_{H_{x}^{4}})\|u_{2}(\tau)\|_{H_{x}^{4}}\|w(\tau)\|_{H_{x}^{3}}d\tau$
$+\int_{t}^{+\infty}(\|u_{1}(\tau)\|_{W_{x}^{4,\infty}}+\|u_{2}(\tau)\|_{H_{x}^{4}})\|w(\tau)\|_{H_{x}^{3}}^{2}d\tau$
$+\int_{t}^{+\infty}(\|u_{1}(\tau)\|_{W_{x}^{3,\infty}}+\|u_{2}(\tau)\|_{H_{x}^{3}})\|u_{2}(\tau)\|_{H_{x}^{3}}\|w(\tau)\|_{H_{x}^{3}}^{2}d\tau$
$+\int_{t}^{+\infty}(\|\partial_{x_{1}}w(\tau)\|_{L_{x}^{\infty}}+\|\partial_{x_{1}}w(\tau)\|_{W_{x}^{1,\frac{2}{\nu}}})\|w(\tau)\|_{H_{x}^{3}}^{2}d\tau$
$+\int_{t}^{+\infty}(\|u_{1}(\tau)\|_{W_{x}^{3,\infty}}+\|u_{2}(\tau)\|_{H_{x}^{3}})\|w(\tau)\|_{H_{x}^{3}}^{3}d\tau$
$+\int_{t}^{+\infty}\|w(\tau)\|_{H_{x}^{3}}^{4}d\tau$
$+\int_{t}^{+\infty}(\|u_{1}(\tau)\|_{W_{x}^{1,\infty}}+\|u_{2}(\tau)\|_{H_{x}^{3}})\|u_{2}(\tau)\|_{H_{x}^{3}}\|w(\tau)\|_{L_{x}^{2}}^{13}d\tau$
$+\int_{t}^{+\infty}(\|u_{1}(\tau)\|_{W_{x}^{1,\infty}}+\|u_{2}(\tau)\|_{H_{x}^{3}})\|w(\tau)\|_{L_{x}^{2}}^{14}d\tau.$
By the Sobolev embedding, we find
$\|\partial_{x_{1}}w\|_{L_{x}^{\infty}}\lesssim\|\partial_{x_{1}}w\|_{W_{x}^{2\nu,\frac{2}{\nu}}}\lesssim\||\partial_{x_{1}}|^{\frac{\nu}{2}}w\|_{W_{x}^{1+\frac{3}{2}\nu,\frac{2}{\nu}}}\lesssim\||\partial_{x_{1}}|^{\frac{\nu}{2}}w\|_{W_{x}^{2,\frac{2}{\nu}}}.$
Hence the Hölder inequality and Lemma 3.3 yield
$\|w(t)\|_{H_{x}^{3}}^{2}+6\int_{\mathbb{R}^{3}}w(\Delta w)^{2}dx+M\|w\|_{L_{x}^{2}}^{14}$
$\lesssim$ $(r^{3}+r^{4})\int_{t}^{+\infty}\tau^{-2-\frac{2}{3}\delta}\|w(\tau)\|_{H_{x}^{3}}d\tau+(r+r^{4})\int_{t}^{+\infty}\tau^{-1-\frac{\delta}{3}}\|w(\tau)\|_{H_{x}^{3}}^{2}d\tau$
$+\|w(\tau)\|_{L_{\tau}^{\frac{6}{3+2\nu}}(t,\infty;H_{x}^{3})}^{2}\||\partial_{x_{1}}|^{\frac{\nu}{2}}w(\tau)\|_{L_{\tau}^{\frac{6}{3-2\nu}}(t,\infty;W_{x}^{2,\frac{2}{\nu}})}$
$+(r+r^{2})\int_{t}^{+\infty}\tau^{-1-\frac{\delta}{3}}\|w(\tau)\|_{H_{x}^{3}}^{3}d\tau+\int_{t}^{+\infty}\|w(\tau)\|_{H_{x}^{3}}^{4}d\tau$
$+(r^{3}+r^{4})\int_{t}^{+\infty}\tau^{-2-\frac{2}{3}\delta}\|w(\tau)\|_{H_{x}^{3}}^{13}d\tau+(r+r^{2})\int_{t}^{+\infty}\tau^{-1-\frac{\delta}{3}}\|w(\tau)\|_{H_{x}^{3}}^{14}d\tau.$
Therefore,
$\|w(t)\|_{H_{x}^{3}}^{2}+6\int_{\mathbb{R}^{3}}w(\Delta w)^{2}dx+M\|w\|_{L_{x}^{2}}^{14}$
$\lesssim$ $(r^{3}+r^{4})t^{-\alpha-\frac{2}{3}\delta-1}\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)$
$+(r+r^{4})t^{-2\alpha-\frac{\delta}{3}}\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)^{2}$
$+t^{-3\alpha+\frac{\nu}{3}+\frac{1}{2}}\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)^{2}$
$\qquad\qquad\qquad\quad\quad\times\left(\sup_{t\in[T,\infty)}t^{\alpha}\||\partial_{x_{1}}|^{\frac{\nu}{2}}w(\tau)\|_{L_{\tau}^{\frac{6}{3-2\nu}}(t,\infty;W_{x}^{2,\frac{2}{\nu}})}\right)$
$+(r+r^{2})t^{-3\alpha-\frac{\delta}{3}}\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)^{3}+t^{-4\alpha+1}\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)^{4}$
$+(r^{3}+r^{4})t^{-13\alpha-\frac{2}{3}\delta-1}\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)^{13}$
$+(r+r^{2})t^{-14\alpha-\frac{\delta}{3}}\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)^{14}$
for any $t\in[T,\infty)$, where the implicit constants are independent of $\lambda$ and $\mu$. By the Gagliardo-Nirenberg inequality
$\|\Delta w\|_{L_{x}^{4}}\lesssim\|w\|_{L_{x}^{2}}^{\frac{1}{12}}\|w\|_{H_{x}^{3}}^{\frac{11}{12}},$
we have
$6\int_{\mathbb{R}^{3}}w(\Delta w)^{2}dx$ $\leq$ $6\|w\|_{L_{x}^{2}}\|\Delta w\|_{L_{x}^{4}}^{2}$
$\leq$ $C\|w\|_{L_{x}^{2}}^{\frac{7}{6}}\|w\|_{H_{x}^{3}}^{\frac{11}{6}}\ \leq C^{\prime}\|w\|_{L_{x}^{2}}^{14}+\frac{1}{2}\|w\|_{H_{x}^{3}}^{2}.$
Thus, if $M>0$ is sufficiently large, then
(4.14) $\|w(t)\|_{H_{x}^{3}}^{2}\sim\|w(t)\|_{H_{x}^{3}}^{2}+6\int_{\mathbb{R}^{3}}w(\Delta w)^{2}dx+M\|w\|_{L_{x}^{2}}^{14}.$
Combining (LABEL:y1) with (4.14), we have (4.1). ∎
Next we derive the estimates for $|\partial_{x_{1}}|^{\frac{\nu}{2}}w$ in $L_{\tau}^{\frac{6}{3-2\nu}}(t,\infty;W_{x}^{2,\frac{2}{\nu}})$.
Lemma 4.2.
Let $w$ be a solution to (4). Then we have
$\sup_{t\in[T,\infty)}t^{\alpha}\||\partial_{x_{1}}|^{\frac{\nu}{2}}w(\tau)\|_{L_{\tau}^{\frac{6}{3-2\nu}}(t,\infty;W_{x}^{2,\frac{\nu}{2}})}$
$\lesssim$ $(r^{3}+r^{4})T^{\alpha-\frac{2}{3}\delta-1}+(r+r^{2})T^{-\frac{\delta}{3}}\|w\|_{Z_{T}}$
$+T^{-\alpha+\frac{\nu}{3}+\frac{1}{2}}\|w\|_{Z_{T}}^{2},$
where the implicit constants are independent of $\lambda$ and $\mu$.
Proof of Lemma 4.2..
Since $w$ satisfies
$w(t)$ $=$ $-(1+\lambda t)^{-5}\rho_{\mu}\ast\partial_{x_{1}}\int_{t}^{\infty}V(t-\tau)\left[w(\tau)^{2}\right]d\tau$
$-2(1+\lambda t)^{-5}\rho_{\mu}\ast\partial_{x_{1}}\int_{t}^{\infty}V(t-\tau)\left[(u_{1}(\tau)+u_{2}(\tau))w(\tau)\right]d\tau$
$-(1+\lambda t)^{-5}\rho_{\mu}\ast\partial_{x_{1}}\int_{t}^{\infty}V(t-\tau)\left[2u_{1}(\tau)u_{2}(\tau)+u_{2}(\tau)^{2}\right]d\tau,$
applying the Strichartz estimates (Lemma 2.1 (2.3)), we have
$\||\partial_{x_{1}}|^{\frac{\nu}{2}}w(\tau)\|_{L_{\tau}^{\frac{6}{3-2\nu}}(t,\infty;W_{x}^{2,\frac{2}{\nu}})}$
$\lesssim$ $\||\partial_{x_{1}}|^{1-\frac{\nu}{2}}w(\tau)^{2}\|_{L_{\tau}^{\frac{6}{3+2\nu}}(t,\infty;W_{x}^{2,\frac{2}{2-\nu}})}+\left\|(u_{1}(\tau)+u_{2}(\tau))w(\tau)\right\|_{L_{\tau}^{1}(t,\infty;H_{x}^{3})}$
$+\|2u_{1}(\tau)u_{2}(\tau)+u_{2}(\tau)^{2}\|_{L_{\tau}^{1}(t,\infty;H_{x}^{3})}.$
By the Hölder inequality,
$\||\partial_{x_{1}}|^{1-\frac{\nu}{2}}w(\tau)^{2}\|_{L_{\tau}^{\frac{6}{3+2\nu}}(t,\infty;W_{x}^{2,\frac{2}{2-\nu}})}$
$\lesssim$ $\|w(\tau)^{2}\|_{L_{\tau}^{\frac{6}{3+2\nu}}(t,\infty;W_{x}^{3,\frac{2}{2-\nu}})}$
$\lesssim$ $\|\|w(\tau)\|_{H_{x}^{3}}^{2}\|_{L_{\tau}^{\frac{6}{3+2\nu}}(t,\infty)}$
$\lesssim$ $\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)^{2}\|\tau^{-2\alpha}\|_{L_{\tau}^{\frac{6}{3+2\nu}}(t,\infty)}$
$\lesssim$ $t^{-2\alpha+\frac{\nu}{3}+\frac{1}{2}}\|w\|_{Z_{T}}^{2}.$
By the Hölder inequality and Lemma 3.3,
$\|(u_{1}(\tau)+u_{2}(\tau))w(\tau)\|_{L^{1}(t,\infty;H_{x}^{3})}$
$\lesssim$ $\left\|(\|u_{1}(\tau)\|_{W_{x}^{3,\infty}}+\|u_{2}(\tau)\|_{H_{x}^{3}})\|w(\tau)\|_{H_{x}^{3}}\right\|_{L_{\tau}^{1}(t,\infty)}$
$\lesssim$ $(r+r^{2})\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)\|\tau^{-1-\alpha-\frac{\delta}{3}}\|_{L_{\tau}^{1}(t,\infty)}$
$\lesssim$ $(r+r^{2})t^{-\alpha-\frac{\delta}{3}}\|w\|_{Z_{T}},$
$\|2u_{1}(\tau)u_{2}(\tau)+u_{2}(\tau)^{2}\|_{L^{1}(t,\infty;H_{x}^{3})}$
$\lesssim$ $\left\|\|u_{1}(\tau)\|_{W_{x}^{3,\infty}}\|u_{2}(\tau)\|_{H_{x}^{3}}+\|u_{2}(\tau)\|_{H_{x}^{3}}^{2}\right\|_{L_{\tau}^{1}(t,\infty)}$
$\lesssim$ $(r^{3}+r^{4})\|\tau^{-2-\frac{2}{3}\delta}\|_{L_{\tau}^{1}(t,\infty)}$
$\lesssim$ $(r^{3}+r^{4})t^{-1-\frac{2}{3}\delta}.$
Substituting (4), (4), (4) into (LABEL:cF), we have (4.2), where the implicit constants are independent of $\lambda$ and $\mu$. ∎
Proof of Theorem 1.1..
By Lemmas 4.1 and 4.2, we have
$\|w\|_{Z_{T}}$ $\lesssim$ $(r^{3}+r^{4})T^{\alpha-\frac{2}{3}\delta-1}+(r+r^{4})T^{-\frac{\delta}{3}}\|w\|_{Z_{T}}+T^{-\alpha+\frac{\nu}{3}+\frac{1}{2}}\|w\|_{Z_{T}}^{2}$
$+(r+r^{2})T^{-\alpha-\frac{\delta}{3}}\|w\|_{Z_{T}}^{2}+T^{-2\alpha+1}\|w\|_{Z_{T}}^{3}$
$+(r^{3}+r^{4})T^{-11\alpha-\frac{2}{3}\delta-1}\|w\|_{Z_{T}}^{12}+(r+r^{2})T^{-12\alpha-\frac{\delta}{3}}\|w\|_{Z_{T}}^{13}.$
We now choose $\alpha,\nu>0$ so that $1/2<\alpha<1$ and $\nu/3+1/2<\alpha$. Then, we see that there exists $T>0$ which depends on $r$ and is independent of $\lambda$ and $\mu$ such that for any $0<\lambda<1$ and $0<\mu<1$,
(4.20) $\|w\|_{Z_{T}}\leq 2r.$
Combining a priori estimate (4.20) with the standard compactness argument (see [22, Section 3] for instance), we find that there exists a unique solution $u\in C([T,\infty);H_{x}^{1}(\mathbb{R}))$ to (1.1) which satisfies $\|w\|_{Z_{T}}=\|u-u_{1}-u_{2}\|_{Z_{T}}\leq 2r$. By conservations of the mass (1.2) and the energy (1.3), we see $u\in C(\mathbb{R};H^{1}(\mathbb{R}))$. Furthermore, from the above inequality and Lemma 3.3, we see
$\|u(t)-V(t)u_{+}\|_{H_{x}^{3}}$ $\lesssim$ $\|w(t)\|_{H_{x}^{3}}+\|u_{2}(t)\|_{H_{x}^{3}}$
$\lesssim$ $rt^{-\alpha}+r^{2}t^{-1-\frac{\delta}{3}}$
$\lesssim$ $(r+r^{2})t^{-\alpha}$
for any $t\geq 1$. This completes the proof of Theorem 1.1. ∎
Appendix A Proof of Lemma 3.2.
In this appendix, we prove Lemma 3.2.
Proof of Lemma 3.2..
By (3.3), we see
(A.1) $\partial_{\eta_{1}}\phi(\xi,\eta)$ $=$ $3\xi_{1}^{2}+\xi_{2}^{2}+\xi_{3}^{2}-6\xi_{1}\eta_{1}-2\xi_{2}\eta_{2}-2\xi_{3}\eta_{3},$
(A.2) $\partial_{\eta_{2}}\phi(\xi,\eta)$ $=$ $2\xi_{1}\xi_{2}-2\xi_{1}\eta_{2}-2\eta_{1}\xi_{2},$
(A.3) $\partial_{\eta_{3}}\phi(\xi,\eta)$ $=$ $2\xi_{1}\xi_{3}-2\xi_{1}\eta_{3}-2\eta_{1}\xi_{3}.$
Hence, we have
$\phi(\xi,\eta)-\eta\cdot\nabla_{\eta}\phi(\xi,\eta)=(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})\xi_{1}+2\eta_{1}(\eta_{2}\xi_{2}+\eta_{3}\xi_{3}).$
Therefore,
(A.4) $\eta_{2}\xi_{2}+\eta_{3}\xi_{3}=-\frac{3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}}{2\eta_{1}}\xi_{1}+N_{1},$
where $N_{1}=N_{1}(\xi,\eta)$ is given by
(A.5) $N_{1}=\frac{1}{2\eta_{1}}\left(\phi(\xi,\eta)-\eta\cdot\nabla_{\eta}\phi(\xi,\eta)\right).$
[…]
$2\eta_{1}(\eta_{3}\xi_{2}-\eta_{2}\xi_{3})$ $=$ $(\xi_{3}-\eta_{3})\partial_{\eta_{2}}\phi-(\xi_{2}-\eta_{2})\partial_{\eta_{3}}\phi.$
Therefore,
(A.6) $\eta_{3}\xi_{2}-\eta_{2}\xi_{3}$ $=$ $\frac{1}{2\eta_{1}}\left\{(\xi_{3}-\eta_{3})\partial_{\eta_{2}}\phi-(\xi_{2}-\eta_{2})\partial_{\eta_{3}}\phi\right\}.$
By (A.4) and (A.6),
$(\eta_{2}^{2}+\eta_{3}^{2})(\xi_{2}^{2}+\xi_{3}^{2})$ $=$ $(\eta_{2}\xi_{2}+\eta_{3}\xi_{3})^{2}+(\eta_{3}\xi_{2}-\eta_{2}\xi_{3})^{2}$
$=$ $\frac{(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{2}}{4\eta_{1}^{2}}\xi_{1}^{2}+(\eta_{2}^{2}+\eta_{3}^{2})N_{2}(\xi,\eta),$
where $N_{2}=N_{2}(\xi,\eta)$ is given by
$N_{2}$ $=$ $-\frac{3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}}{2\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\xi_{1}\left(\phi-\eta\cdot\nabla_{\eta}\phi\right)+\frac{1}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\left(\phi-\eta\cdot\nabla_{\eta}\phi\right)^{2}$
$+\frac{1}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\left\{(\xi_{3}-\eta_{3})\partial_{\eta_{2}}\phi-(\xi_{2}-\eta_{2})\partial_{\eta_{3}}\phi\right\}^{2}$
$=$ $\frac{1}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\{-(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})\xi_{1}+2\eta_{1}\eta_{2}\xi_{2}+2\eta_{1}\eta_{3}\xi_{3}\}\phi$
$+\frac{1}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\{(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})\xi_{1}-2\eta_{1}\eta_{2}\xi_{2}-2\eta_{1}\eta_{3}\xi_{3}\}\eta_{1}\partial_{\eta_{1}}\phi$
$+\frac{1}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\{2\eta_{1}\eta_{3}\xi_{2}\xi_{3}-2\eta_{1}\eta_{2}\xi_{3}^{2}+(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})\eta_{2}\xi_{1}$
$\qquad\qquad\qquad\ -2\eta_{1}(\eta_{2}^{2}+\eta_{3}^{2})\xi_{2}\}\partial_{\eta_{2}}\phi$
$+\frac{1}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\{-2\eta_{1}\eta_{3}\xi_{2}^{2}+2\eta_{1}\eta_{2}\xi_{2}\xi_{3}+(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})\eta_{3}\xi_{1}$
$\qquad\qquad\qquad\ -2\eta_{1}(\eta_{2}^{2}+\eta_{3}^{2})\xi_{3}\}\partial_{\eta_{3}}\phi.$
Hence
(A.8) $\xi_{2}^{2}+\xi_{3}^{2}=\frac{(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{2}}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\xi_{1}^{2}+N_{2}(\xi,\eta).$
Substituting (A.4) and (A.8) into (A.1), we have
$\partial_{\eta_{1}}\phi(\xi,\eta)$
$=$ $(3\xi_{1}^{2}-6\xi_{1}\eta_{1})+(\xi_{2}^{2}+\xi_{3}^{2})-2(\eta_{2}\xi_{2}+\eta_{3}\xi_{3})$
$=$ $(3\xi_{1}^{2}-6\xi_{1}\eta_{1})+\frac{(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{2}}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\xi_{1}^{2}+\frac{3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}}{\eta_{1}}\xi_{1}-2N_{1}+N_{2}$
$=$ $\frac{9\eta_{1}^{4}+18\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+(\eta_{2}^{2}+\eta_{3}^{2})^{2}}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\xi_{1}^{2}-\frac{3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2}}{\eta_{1}}\xi_{1}-2N_{1}+N_{2}.$
Therefore
$\frac{9\eta_{1}^{4}+18\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+(\eta_{2}^{2}+\eta_{3}^{2})^{2}}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\xi_{1}^{2}-\frac{3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2}}{\eta_{1}}\xi_{1}$
$=$ $\partial_{\eta_{1}}\phi+2N_{1}-N_{2}.$
[…]
$2(\xi_{1}-\eta_{1})(\eta_{2}\xi_{2}+\eta_{3}\xi_{3})-2\xi_{1}(\eta_{2}^{2}+\eta_{3}^{2})=\eta_{2}\partial_{\eta_{2}}\phi+\eta_{3}\partial_{\eta_{3}}\phi.$
Hence by (A.4), we obtain
$2(\xi_{1}-\eta_{1})\left\{-\frac{3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}}{2\eta_{1}}\xi_{1}+N_{1}\right\}-2\xi_{1}(\eta_{2}^{2}+\eta_{3}^{2})=\eta_{2}\partial_{\eta_{2}}\phi+\eta_{3}\partial_{\eta_{3}}\phi.$
Therefore,
$\frac{3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}}{\eta_{1}^{2}}\xi_{1}^{2}-\frac{3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2}}{\eta_{1}}\xi_{1}$
$=$ $\frac{1}{\eta_{1}}\left\{2(\xi_{1}-\eta_{1})N_{1}-\eta_{2}\partial_{\eta_{2}}\phi-\eta_{3}\partial_{\eta_{3}}\phi\right\}.$
By ${(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})/\eta_{1}^{2}\times}$ $\{9\eta_{1}^{4}+18\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+(\eta_{2}^{2}+\eta_{3}^{2})^{2}\}/\{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})\}\times(A)$, we have
$3\frac{(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{2}}{4\eta_{1}^{3}(\eta_{2}^{2}+\eta_{3}^{2})}\xi_{1}=N_{3},$
where $N_{3}=N_{3}(\xi,\eta)$ is given by
$N_{3}$ $=$ $\frac{3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}}{\eta_{1}^{2}}(\partial_{\eta_{1}}\phi+2N_{1}-N_{2})$
$-\frac{9\eta_{1}^{4}+18\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+(\eta_{2}^{2}+\eta_{3}^{2})^{2}}{4\eta_{1}^{3}(\eta_{2}^{2}+\eta_{3}^{2})}\{2(\xi_{1}-\eta_{1})N_{1}-\eta_{2}\partial_{\eta_{2}}\phi-\eta_{3}\partial_{\eta_{3}}\phi\}.$
Hence we obtain
(A.12) $\xi_{1}=\frac{4\eta_{1}^{3}(\eta_{2}^{2}+\eta_{3}^{2})}{3(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{2}}N_{3}.$
Substituting (A.5), (LABEL:n2), (LABEL:n3) into (A.12), we have (3.5), where
$A_{0}(\eta)$ $=$ $\frac{1}{3p(\eta)}\{9\eta_{1}^{4}+30\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+5(\eta_{2}^{2}+\eta_{3}^{2})^{2}\},$
$A_{1}(\eta)$ $=$ $-\frac{1}{3p(\eta)}\eta_{1}\{9\eta_{1}^{4}+18\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+(\eta_{2}^{2}+\eta_{3}^{2})^{2}\},$
$A_{j}(\eta)$ $=$ $-\frac{4}{3p(\eta)}\eta_{i}(\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad j=2,3,$
$B_{0,1}(\eta)$ $=$ $-\frac{4}{p(\eta)}\eta_{1}(\eta_{2}^{2}+\eta_{3}^{2}),$
$B_{0,k}(\eta)$ $=$ $-\frac{2}{3p(\eta)}\eta_{k}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad k=2,3,$
$B_{1,1}(\eta)$ $=$ $\frac{4}{p(\eta)}\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2}),$
$B_{1,k}(\eta)$ $=$ $\frac{2}{3p(\eta)}\eta_{1}\eta_{k}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad k=2,3,$
$B_{j,1}(\eta)$ $=$ $\frac{4}{p(\eta)}\eta_{1}\eta_{j}(\eta_{2}^{2}+\eta_{3}^{2}),\quad j=2,3,$
$B_{j,2}(\eta)$ $=$ $\frac{2}{3p(\eta)}(\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad j=2,3,$
$C_{j,1}(\eta)$ $=$ $-\frac{2}{3p(\eta)}\eta_{5-j}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad j=2,3,$
$C_{j,2}(\eta)$ $=$ $\frac{2}{3p(\eta)}\eta_{j}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad j=2,3,$
with $p(\eta)=(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{2}$. Hence we have Lemma 3.2. ∎
Acknowledgements
J.S was supported by JSPS KAKENHI Grant Numbers JP25H00597 and JP23K20805.